# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 5 Assurance and Alignment for AGI
In our opinion, the only feasible means for anticipation and control of potential disruptions due to AI is self-imposed or mandatory scrutiny within the organizations developing the technology [^33] As precedent, the FAA has Designated Engineering Representatives (DERs) within the aviation industry: the argument being that only those actually working on the products know enough of the details to identify and anticipate hazards. This is obviously vulnerable to regulatory capture, as demonstrated by the Boeing MCAS scandal, but is now reinforced and has otherwise worked well. combined with public disclosure and debate. We believe there is a strong role for modern assurance methods within this framework, rather than the “codes of practice” envisioned in the EU regulations, especially for systems that perform specific functions. The reason that other industries have moved from standards and codes of practice to assurance cases is hard-won experience that assurance has to be based on identification and elimination of the hazards of the specific system and environment under consideration. For general-purpose AFGI systems, such as advanced LLMs, we believe that mechanistic protections such as guards and internal monitors may prove inadequate and that new methods will need to be developed based on improved understanding of topics such as emergent behavior [101], the cognitive basis of language and shared intentionality (teamwork) [167, 168, 20, 149], and the sociology of judgment and cooperation [79].
@@
So far, our discussion has considered AFGI systems that are not far beyond those already current, and has not touched on “real” AGI. We suspect that popular opinion misinterprets advanced AFGI as AGI and thereby underestimates the potential danger of true AGI. As widely quoted (and variously attributed) “it is difficult to make predictions, especially about the future,” and we consider the technology and the timing of true AGI to be unknowable. [^34] But see the One Hundred Year Study on Artificial Intelligence https://ai100.stanford.edu/. Hence, it is impossible to identify potential hazards and suitable mitigations and methods of assurance. However, it does seem sensible to consider possible scenarios, even speculative ones (so science fiction may be an appropriate forum to perform this work). For example, as already noted, one dystopian possibility is that AGI may misinterpret its goals or establish its own at variance to our desires and best interests, or that we may simply give it poorly chosen goals. Hadfield-Menell and colleagues [67] provide examples of poorly specified goals (e.g., King Midas’ wish that everything he touches turn to gold) or misinterpreted ones (e.g., a robot given the goal of cleaning up dirt repeatedly dumps and cleans the same dirt). They identify the generic problem as one of value alignment and propose “Cooperative Inverse Reinforcement Learning” (CIRL) as a promising technical framework, later generalized to “assistance games” [151]. The idea is that the AGI’s goal is continuously to learn what it is we want it to do by observing our behavior (but what if its “master” has bad intent?). A more wide-ranging and speculative proposal is for “Guaranteed Safe AI” [46].
@@
Beyond AGI is the concern that machines may become conscious [150]. This is a difficult topic because there is no agreed definition for consciousness, nor ways of detecting or measuring it (consider debate over whether animals are conscious). And AI consciousness may exist but be utterly unlike our own (for an analogy, consider the possibility of consciousness distributed among the nine “brains” of an octopus [38, 118]). Nonetheless, there is general agreement on two aspects of consciousness: intentional consciousness is the ability to direct attention and to think about something and to know that you are doing so, while phenomenal consciousness is “what it’s like” to have subjective experiences such as the smell of a rose or the feeling of pain. There are two subtopics here: one is how plausible or likely it is that AI could achieve either kind of consciousness, and the second is what the consequences might be.
@@
Having said that, there is no agreement how human consciousness is achieved, nor what purpose it serves. However, there have been experimental attempts to construct machine consciousness [140, 61]. Mostly these are based on the idea that consciousness derives from explicit models of self and of others, and is related to communication and language [74, 130] (these ideas correspond most closely with Higher Order Thought, or HOT, models of human consciousness [33, 63, 149]). None of these experiments have delivered any sign of consciousness, although they have sharpened some of the questions, and the possibility of machine consciousness remains open.
Were it to be achieved, intentional AI consciousness seems unlikely to add capabilities beyond those of AGI: intentional consciousness seems to be essential to human reasoning, but AI produces facsimiles of reasoning by other means, just as airplanes fly without growing feathers or flapping their wings. However, AI consciousness might require us to attribute some of the system’s behavior to conscious decisions, thereby raising philosophically difficult questions regarding moral responsibility, personal identity, and free will.
Phenomenal consciousness would also raise philosophically difficult questions, but in this case they would concern ethics—not what AI might do to us, but how we should treat AI. An AI system that truly has subjective experience, that feels pain and pleasure, raises profound questions on the foundations of ethics and on just treatment of other sentient beings. It also asks how we could tell whether the AI’s subjective experience is real or faked (cf. “philosophical zombies” [94]).
@@ 6 Summary and Conclusion
We have described and discussed methods to ensure and assure critical properties of AI systems from the traditional “dependability” perspective. This requires that assurance for critical properties rests only on elements of the system for which we have near-complete understanding of what they do, how they do it, why they do it, and the environment in which they do it, and comparably complete understanding of their implementation and its correctness. All of these topics should be addressed in an assurance case that assembles claims, evidence, arguments, and theories in a manner that provides sufficient confidence in its top (i.e., overall) claim to justify deployment.
@@
The dependability perspective asserts that systems based on AI and Machine Learning (ML) cannot satisfy these requirements; the contrary “trustworthy” perspective believes they can, in some cases. There is a continuum or spectrum between these viewpoints and all are worthy of investigation; we focus here on those at the dependability end and hope others will survey other points along the spectrum.
@@
In that regard, we note recent papers that propose to apply assurance or safety cases to the trustworthy perspective (e.g., [44, 46]). We welcome these developments but stipulate that when we speak of an assurance case we set a high bar. In particular, the case must be indefeasible, meaning there is no credible new information that would change its assessment [148]. A (deliberate) consequence of this is that the case must have “no gaps,” which generally requires the argument to be deductive [29]; exceptions must be noted as residual doubts and shown to pose negligible risk. Evidence must be assessed skeptically (we advocate use of confirmation measures [64]) and shown to support not merely a claim of something measured (e.g., “we did penetration testing to such and such standard”) but one of something useful (e.g., “therefore faults of type xx are absent”). In well-developed technical areas, the assurance argument should largely be assembled from standard theories that provide well-attested (and ideally “pre-certified”) subcases for common fragments of the overall case (e.g., static analysis for absence of certain coding faults). The upper levels of the assurance case do not concern the design or implementation of the system, but a description of its environment and assumptions, derivation of its hazards, and development of its requirements, all bound together by dependability requirements validation to demonstrate that the requirements mitigate the hazards. Next comes the specification that describes the design of the system, coupled with intent verification to show that these satisfy the requirements. Only then do we get to the implementation of the system and its correctness verification where, we stress, for critical systems we need near certainty that this satisfies the specification. We suggest that formal verification of properties of neural networks often corresponds more closely to static analysis that full correctness verification. Due to these challenging stipulations, we do not consider that assurance for trustworthy AI behavior is feasible in most current applications, although we welcome continued research in these directions and look forward to positive developments. We also acknowledge that many AI applications do not require the same degree of assurance as safety critical cyber-physical systems (CPS), but it is challenging to determine which parts of a full assurance case may then be relaxed, and how, and by how much (probabilistic estimation may be useful [29, Section3]).
@@
Work by Dong and colleagues [49] meets many of our stipulations but much of the evidence for their assurance case comes from statistical modeling of the reliability of its ML components. This is reasonable from a trustworthiness perspective but less so from the dependability viewpoint, where we would want to see runtime verification and dynamic assurance monitoring. The dependability viewpoint does not trust AI and ML elements because they are derived by experimental optimization (“training”) and their complete behavior is unknown. Consequently, assurance is achieved using runtime verification with guarded architectures in which overall behavior is checked for required properties by traditionally developed and assured guards. An obstacle to this approach is that the guards must accurately perceive (i.e., build a model of) the current state of the system’s environment, and this may require AI and ML (e.g., for a self-driving car to perceive other road users).
@@
Next, we looked at foundation models such as LLMs and other general-purpose systems on which more specific systems can be constructed. The additional problem here is that the ultimate application and its hazards are unknown and so the general-purpose system has to use very general guards. We considered guards based on normative principles, such as ethics and other overarching frameworks, and we also examined explanations and reputation systems. Again, we found that effective guards may be feasible, but the main difficulty is assurance that the world view constructed by the guard aligns with human perception.
@@
Finally, we considered AGI and other futuristic prospects such as the singularity and machine consciousness. Here, we are concerned less about faults and failures of the system and more about unanticipated and disruptive consequences of their correct and intended behavior. It is infeasible to guard these systems so we briefly examined ways to ensure that their goals align with ours (which, again, is largely a question of how the world is perceived). Although we concede that near-term development of AGI is possible, we suggest that a more imminent concern is AFGI (Artificial Fairly Good Intelligence), which can be projected from current advances in generative AI such as LLMs, and we identified some proposed government regulations and other safeguards.
@@
Overall, we suggest that for CPS and other systems for specific functions, architectures similar to those labeled 4, 7 and 8 in Sections 2.4 and 2.5 are (in ascending order) the most attractive. These use a traditionally engineered and assured guard for “last second” protection (4), plus weakly assured but diverse perception for defense in depth (7), and assured detection of micro-ODDs with specialized guards for each one (8). Similar architectures can be adapted for general-purpose systems, but with less assurance (as the hazards are less well-defined and perception more difficult). For AFGI, and even more so for AGI, we suggest that protection and alignment must go beyond restrictions on their mechanisms and build on an understanding of the cognitive basis of language and intelligence and the sociology of group cooperation.
There are three research topics that we consider urgent;
@@
    Much current work focuses on development of guards, but these are dependent on accurate perception of the world and we consider that assured perception is a neglected topic. This applies not only to explicit models (e.g., location of other vehicles) as constructed by CPS, but implicit ones (e.g., current social context and role within it) underlying the behavior of LLMs.
    Assured perception systems need not provide fully detailed world models, but must be accurate to the level of resolution or discrimination needed for modest but effective guards (e.g., detection of micro-ODDs and prediction of urgent actions). Research from the trustworthiness perspective could usefully be applied to this problem. Related to assured perception is the topic of principled fusion of diverse world models, possibly having different levels of resolution (e.g., one fully detailed, and another less detailed but assuredly accurate). A promising approach uses “predictive processing” and a “dual process” cognitive structure similar to the human brain.
  • Development of layered, recursively structured architectures and associated guards to provide defense in depth, within a socio-technical approach to design (i.e., one that integrates technology with human and community aspects).
    It is unlikely that a single guard can provide full protection for an AI system in a complex environment, so overall architectures should favor defense in depth based on rational principles, which begin with careful appraisal of what the system is to do, and the environment in which it will operate. In particular, environments where perception is easier should be favored (as with ODDs for self-driving cars) and possibly constructed (e.g., dedicated tracks for an automated shuttle bus) and those with “tight coupling” and “interactive complexity” [129] should be avoided. In addition to architectures and to design and assurance methods for strong guards, research should explore credible methods for dynamic assurance and resilience for those applications where strongly assured guards are infeasible.
  • Research on the computational mechanisms underlying emergent behavior and the cognitive basis of language, shared intentionality, and intelligence.
    We cannot develop methods for assurance and alignment of AFGI and AGI without better understanding of the computational processes underlying cognition, and human (and to some extent animal) cognition are the only models that we have. Beyond cognition, we need to understand the processes underlying teamwork (i.e., shared intentionality) and the sociology of cooperative behavior.

### v3

@@ 5.2 Resilience as a Key Response to AFGI/AGI
In our opinion, a vital means for anticipation and control of potential disruptions due to AI is self-imposed or mandatory scrutiny within the organizations developing the technology, [^43] As precedent, the FAA has Designated Engineering Representatives (DERs) within the aviation industry: the argument being that only those actually working on the products know enough of the details to identify and anticipate hazards. This is obviously vulnerable to regulatory capture, as demonstrated by the Boeing MCAS scandal, but is now reinforced and has otherwise worked well. combined with public disclosure and debate leading to establishment of regulatory goals. We believe there is a strong role for modern methods such as assurance cases and goal-based regulation within this framework, rather than the “codes of practice” envisioned in the EU regulations, especially for systems that perform specific functions. The reason that other industries have moved from standards and codes of practice to explicit goals and assurance cases is hard-won experience that assurance has to be based on identification and elimination of the hazards of the specific system and environment under consideration. For failures precipitated by AFGI systems, such as advanced LLMs, we believe that mechanistic protections such as guards and internal monitors may prove inadequate and that new methods will need to be developed based on improved understanding of topics such as system failure, emergent behavior [143], the cognitive basis of language and shared intentionality (see below), and the sociology of judgment and cooperation [111].
5.3 True AGI
So far, our discussion has considered AFGI systems that are not far beyond those already current, and has not touched on “true” AGI. We suspect that popular opinion misinterprets advanced AFGI as AGI and thereby underestimates how far we are from true AGI and its potential dangers. We have already stated our opinion that true AGI requires more than the facility with language exhibited by AFGI: it requires the ability to think [79], which we interpret as construction and manipulation of general models of the world, and this is not yet in sight. And there is something apart from General Intelligence that makes humans the masters of planet Earth. A single human is a puny thing and no threat to plants or animals, but collectively we wreak our will. And it is not just force that is multiplied by our teamwork: we also increase our collective intelligence. Individually, we do not develop pottery, writing, mathematics—those are products of collective culture. Computers communicate and can be programmed to act collectively, but this is not the same as cultivating alliances, sharing knowledge, “selling” an idea, and ultimately forming a team with shared goals and coordinated plans: that is shared intentionality [231, 232, 27, 209] and it is unique to humans. [^44] Social animals like ants, bees, wolves, dolphins exhibit collective behavior, but this is “programmed in”: a wolf cannot cause its pack to adopt a new idea. Thus, until AGI achieves shared intentionality, it is no great threat to humanity by itself (although in the hands of bad actors it could be): we can always act collectively to outwit it. Of course, this does depend on our willingness to act collectively and history is replete with instances of societies that thought they could exploit a dangerous entity, yet retain control (e.g., Germany in 1933).
@@ 5.3 True AGI
As widely quoted (and variously attributed) “it is difficult to make predictions, especially about the future,” and we consider the technology and the timing of true AGI to be unknowable. [^45] But see the One Hundred Year Study on Artificial Intelligence https://ai100.stanford.edu/. Hence, it is impossible to identify potential hazards and suitable mitigations and methods of assurance. However, it does seem sensible to consider possible scenarios, even speculative ones (so science fiction may be an appropriate forum to perform this work). For example, as already noted, one dystopian possibility is that AGI may misinterpret its goals or establish its own at variance to our desires and best interests, or that we may simply give it poorly chosen goals. Hadfield-Menell and colleagues [95] provide examples of poorly specified goals (e.g., King Midas’ wish that everything he touches turn to gold) or misinterpreted ones (e.g., a robot given the goal of cleaning up dirt repeatedly dumps and cleans the same dirt). They identify the generic problem as one of value alignment and propose “Cooperative Inverse Reinforcement Learning” (CIRL) as a promising technical framework, later generalized to “assistance games” [211]. The idea is that the AGI’s goal is continuously to learn what it is we want it to do by observing our behavior (but what if its “master” has bad intent?). A more wide-ranging and speculative proposal is for “Guaranteed Safe AI” [66], which we outline and criticize in [43, pages 7, 8].
@@
Beyond AGI is the concern that machines may become conscious [210]. This is a difficult topic because there is no agreed definition for consciousness, nor ways of detecting or measuring it (consider debate over whether animals are conscious). And AI consciousness may exist but be utterly unlike our own (for an analogy, consider the possibility of a consciousness distributed among the nine “brains” of an octopus [54, 170]). Nonetheless, there is general agreement on two aspects of consciousness: intentional consciousness [^47] The term is a translation from German and should not be considered to focus on “intentions.” is the ability to direct attention and to think about something and to know that you are doing so, while phenomenal consciousness is “what it’s like” to have subjective experiences such as the smell of a rose or the feeling of pain. There are two subtopics here: one is how plausible or likely it is that AI could achieve either kind of consciousness, and the second is what the consequences might be.
@@
Having said that, there is no agreement how human consciousness is achieved, nor what purpose it serves: Kuhn describes a “landscape” with more than 200 theories of consciousness [145]. However, there have been experimental attempts to construct machine consciousness [198, 85]. Mostly, these are based on the idea that consciousness derives from explicit models of self and of others, and is related to communication and language [104, 186]: these ideas correspond most closely with Higher Order Thought (HOT) models of human consciousness [48, 88, 209]. None of these experiments have delivered any sign of consciousness, although they have sharpened some of the questions, and the possibility of machine consciousness remains open.
Were it to be achieved, intentional AI consciousness seems unlikely to add capabilities beyond those of AGI: intentional consciousness seems to be essential to human reasoning, but AI produces facsimiles of reasoning by other means, just as airplanes fly without growing feathers or flapping their wings. However, AI intentional consciousness might require us to attribute some of the system’s behavior to conscious decisions, thereby raising philosophically difficult questions regarding moral responsibility, personal identity, and free will.
Phenomenal consciousness would also raise philosophically difficult questions, but in this case they would concern ethics—not what AI might do to us, but how we should treat AI. An AI system that truly has subjective experience, that feels pain and pleasure, raises profound questions on the foundations of ethics and on just treatment of other sentient beings [53, 28]. [^48] Several researchers have signed open letters urging caution https://amcs-community.org/open-letters/ and proposing five principles https://conscium.com/open-letter-guiding-research-into-machine-consciousness/ regarding research and development of potentially conscious AI entities. It also asks how we could tell whether the AI’s subjective experience is real or faked (cf. “philosophical zombies” [132]).
@@ 6 Summary and Conclusion
We have described and discussed methods to ensure and assure critical properties of AI systems from the traditional “dependability” perspective. This requires that assurance for critical properties rests only on elements of the system (which may be “guards” rather than operational functions) for which we have near-complete understanding of what they do, how they do it, why they do it, and the environment in which they do it, and comparably complete understanding of their implementation and its correctness. This understanding is needed because testing and operational experience on their own are inadequate for strong assurance: they need to be buttressed by prior confidence in the quality of the system and arguments that the future product and environment will behave like the past. All of these topics should be addressed in an assurance case that assembles claims, evidence, arguments, and theories in a manner that provides sufficient confidence in its top (i.e., overall) claim to justify deployment.
@@
The dependability perspective asserts that systems based on AI and Machine Learning (ML) cannot satisfy the requirements for assurance based on detailed understanding because their inner working is opaque; the contrary “trustworthy” perspective believes they can, in some cases. There is a continuum or spectrum between these viewpoints and all are worthy of investigation; we focus here on those at the dependability end and hope others will survey other points along the spectrum.
@@
In that regard, we note recent papers that propose to apply assurance or safety cases to the trustworthy perspective (e.g., [62, 66]). We welcome these developments but stipulate that when we speak of an assurance case we set a high bar [39]. In particular, the case must be indefeasible, meaning there is no credible new information that would change its assessment [208] and this should be probed skeptically by exploring potential defeaters [38]. A (deliberate) consequence is that the case must have “no gaps,” which generally requires the argument to be deductive [40, 41]; exceptions must be noted as residual doubts and shown to pose negligible risk. [^49] Koopman [138, Section 7.2] asserts that assurance cases for complex systems such as self-driving cars cannot be deductive because they operate in an open world that cannot be fully predicted. We respond that they can still be deductive but some of their elements may be incompletely characterized. The difference is between forcing consideration of weaknesses and doing something about them, such as monitoring at runtime or accepting them as residual risks (i.e., “known unknowns”) versus allowing the argument to have unconsidered gaps (i.e., “unknown unknowns”). Evidence must be assessed skeptically (we advocate use of confirmation measures [89, 206]) and shown to support not merely a claim of something measured (e.g., “we did penetration testing to such and such standard”) but one of something useful (e.g., “therefore faults of type xx are absent”). In well-developed technical areas, the assurance argument should largely be assembled from standard theories that provide well-attested (and ideally “pre-certified”) subcases for common fragments of the overall case (e.g., static analysis for absence of certain coding faults).
The upper levels of the assurance case do not concern the design or implementation of the system, but a description of its environment and assumptions, development of its requirements, and derivation of its hazards, all bound together by dependability requirements validation to demonstrate that the requirements mitigate the hazards. Next comes the specification that describes the design of the system, coupled with intent verification to show that these satisfy the requirements. Only then do we get to the implementation of the system and its correctness verification where, we stress, for critical systems we need near certainty that this satisfies the specification. We suggest that formal verification of properties of neural networks often corresponds more closely to static analysis that full correctness verification.
Due to these challenging stipulations, we do not consider that assurance for trustworthy AI behavior is feasible in most current applications, although we welcome continued research in these directions and look forward to positive developments [43]. We also acknowledge that many AI applications do not require the same degree of assurance as safety-critical cyber-physical systems (CPS), but it is challenging to determine which parts of a full assurance case may then be relaxed, and how, and by how much (probabilistic estimation may be useful [40, Section3]). Work by Dong and colleagues [70] meets many of our stipulations but much of the evidence for their assurance case comes from statistical modeling of the reliability of its ML components. This is reasonable from a trustworthiness perspective but less so from the dependability viewpoint, where we would want to see runtime verification and dynamic assurance monitoring.
@@
The dependability viewpoint does not trust AI and ML elements because they are derived by experimental optimization (“training”) and their complete behavior is unknown. Consequently, assurance is achieved using runtime verification with guarded architectures in which overall behavior is checked for required properties by traditionally developed and assured guards. An obstacle to this approach is that the guards must accurately perceive (i.e., build a model of) the current state of the system’s environment, and this itself may require AI and ML (e.g., for a self-driving car to perceive other road users).
@@
We then considered the idea of system failures (or “normal accidents”) which are failures that are not (primarily) due to component faults but to unanticipated interactions among correct components and subsystems. Because AI systems have facility with language and other modes of human dialog, they can become more deeply embedded into their human social context than intended or recognized. This means that the system boundary is not well defined and makes system failures more likely.
@@
Next, we looked at foundation models such as LLMs and other general-purpose systems on which more specific systems can be constructed (previously these required custom AI components). We acknowledged the possibility of system failures but focused on techniques for guarding general-purpose AI and ML mechanisms. Here, the additional problem is that the ultimate application and its hazards are unknown and so the general-purpose system has to use very general guards. We considered guards based on normative principles, such as ethics and other overarching frameworks, and we also examined explanations and reputation systems. Again, we found that effective guards may be feasible, but the main difficulty is assurance that the world view constructed by the guard aligns with human perception.
@@
Finally, we considered AGI and other futuristic prospects such as the singularity and machine consciousness. Here, the concern is less about faults and failures of AI systems and more about system failures and other unanticipated and disruptive consequences of their correct and intended behavior. It is infeasible to guard these systems so we briefly examined ways to ensure that their goals align with ours (which, again, is largely a question of how the world is perceived). Although we concede that near-term development of AGI is possible, we suggest that a more imminent concern is AFGI (Artificial Fairly General Intelligence), which can be projected from current advances in generative AI such as “frontier” LLMs, and we identified some proposed government regulations and other safeguards.
Many of these are concerned with “existential” risks that we consider inadequately defined: we would like to see far more detailed hazard analysis and more granular (as opposed to “all at once”) scenarios for their emergence. We note that current methods of investigation and assurance (basically “red-teaming”) are woefully incommensurate with the claimed seriousness of the risk. [^50] A recent evaluation reported that the best performing system (OpenAI o1) failed 26% of tests while the worst (DeepSeek R1) failed 100%: https://blogs.cisco.com/security/evaluating-security-risk-in-deepseek-and-other-frontier-reasoning-models.
The acceptable failure rates in any critical system (i.e., one that poses serious risks) are many orders of magnitude more demanding than this.
Proposed mitigations are comparably weak, such as not disclosing the weights [180] in the neural nets of LLMs considered to pose “significantly higher risk”  [^51] Anthropic has a four-level scale for AI Safety Levels (ASLs): ASL-1: smaller models, ASL-2: present large models, ASL-3: significantly higher risk, ASL-4: speculative. [5]; distillation can largely copy the capabilities of an LLM without access to its weights. Hence our call for more precise hazard analysis followed by rational determination of the claims that need to be assured and the confidence required in their assurance. These will require new or revised hazard and failure analysis methods that address the unique aspects of AI and ML (and AFGI in particular): for example, and as mentioned earlier, augmented guidewords for those hazard analysis methods that use them [65, 171].
These methods should address hazards and mitigations over the full development lifecycle, from those focused on growth of foundation models, through the construction of applications built around them, to customers and general society that are impacted by their use. We recommend that mitigations and assurance should focus on the overall system, not just the AI mechanism, and this will extend into human society, where resilience to grave risks should be developed alongside measures for their prevention.
@@
Overall, we suggest that for CPS and other systems for specific functions, architectures similar to those labeled 4, 7 and 8 in Sections 2.5 and 2.6 are (in ascending order) the most attractive. These use a traditionally engineered and assured guard for “last second” protection (4), plus weakly assured but diverse perception for defense in depth (7), and assured detection of micro-ODDs with specialized guards for each one (8). Similar architectures can be adapted for general-purpose systems, but with less assurance (as the hazards are less well-defined and perception more difficult). For AFGI, and even more so for AGI, we suggest that protection and alignment must go beyond restrictions on their mechanisms and build on an understanding of system failures and emergence, the cognitive basis of language and intelligence, and the sociology of group cooperation.
6.1 Research Agenda
There are four research topics that we consider urgent.
  • Methods of hazard analysis are needed that can identify potential sources of failure and harm in complex AI systems, together with design and assurance methods to avoid such harms. Hazard analysis needs to consider the new harms and failure modes of AI systems, their propensity to embed themselves deeply into their social context and thereby widen the system boundary, potentially leading to self-organizing criticality, and system failure (“normal accidents”).
    In addition to hazards, research should examine the credible severity of risks due to AI and ML and should consider chronic and creeping harms as well as catastrophic ones. Research should also distinguish those risks whose mitigation primarily requires technological measures and regulation from those that also require societal response and adaptation.
  • Development of layered, recursively structured architectures and associated guards to provide runtime verification within a socio-technical approach to design (i.e., one that integrates technology with human and community aspects) that avoids both component and system failures.
    It is unlikely that guards alone can provide full protection for an AI system in a complex environment, so overall architectures should favor defense in depth based on diversity and other rational principles, which begin with careful appraisal of what the system is to do, and the environment in which it will operate.
    In addition to architectures and to design and assurance methods that aim to avoid failure, research should explore credible methods for dynamic assurance and post-failure resilience of systems and their social context.
@@ 6.1 Research Agenda
    Much current work focuses on development of guards, but these are dependent on accurate perception of the world and we consider that assured perception is a neglected topic. This applies not only to explicit local models (e.g., location of other vehicles) as constructed by CPS, but implicit ones (e.g., current social context and role within it) underlying the behavior of LLMs.
    Assured perception systems need not provide fully detailed local models, but must be accurate to the level of resolution or discrimination needed for modest but effective guards (e.g., detection of micro-ODDs and prediction of urgent actions). Research from the trustworthiness perspective could usefully be applied to this problem. Related to assured perception is the topic of principled fusion of diverse local models, possibly having different levels of resolution (e.g., one fully detailed, and another less detailed but assuredly accurate). A promising approach uses “predictive processing” and a “dual process” cognitive structure similar to the human brain.
  • Research on the computational mechanisms underlying emergent behavior and the cognitive basis of language, intelligence and shared intentionality,
    We cannot develop methods for assurance and alignment of AFGI and AGI without better understanding of the computational processes underlying cognition including the development and use of domain models. Human (and to some extent animal) cognition are the only models that we have, so techniques such as use of metaphors to map one domain (e.g., social hierarchy) onto another more automated one (e.g., height) [149] may repay study. Beyond cognition, we need to understand the processes underlying teamwork (i.e., shared intentionality) and the sociology of cooperative behavior.